o37. House Numbers A home improvement store has the following house numbers: 335666 left-after a sale. How many different six-digit house numbers can be made from those numbers?
media%2Fa60%2Fa605ca02-3e1b-40fa-83b7-c3
Answer #1

Ans:

Probability of success,p=0.16

n=5

P(x=k)=5Ck*0.16k*(1-0.16)5-k

a)

P(none)=P(x=0)=(1-0.16)^5=0.4182

b)

P(atleast one)=P(x>=1)=1-P(none)=1-0.4182=0.5818

c)

P(all)=P(x=5)=0.16^5=0.0001